Ethanol can undergo decomposition to form two sets of products
C2H5OH----
C2H4(g) + H2O(g)
energy evolved = 45 units/ mole of C2H4
2- CH3CHO(g) + H2(g)
E. evolved = 70 units/ mole of CH3CHO


If the mole ratio of C2H4
 to CH3CHO is 9 : 1 in product gases then the energy involved in decomposition of 1mole ethanol is

 
 

Grade:12th pass

1 Answers

Sunil Kumar FP
9 years ago
we have two sets of reaction
(1)C2H5OH-------C2H4 +H2O
(2)C2H5OH------CH3CHO +H2
Energy involved in the decomposition of ethanol= energy involved in 1st reaction +energy involved in second reaction
energy for first reaction=
AS in the product 9 moles of C2H4 are formed therefore mole of C2H5OH=1/9 mole as the ratio is 9:1
number of mole left=1-1/9=8/9 mole
total energy=energy for 1/9 mole+energy for 8/9 mole
=1/9*45 +8/9*70 unit
=5+62.2 unit
=67.2 unit